OpenStudy (anonymous):

Find the largest natural number n, such that 9999...999 (999 times) is divisible by 9^n

ganeshie8 (ganeshie8):

\(9999\ldots (999 \text{times}) = 10^{999}-1\)

ganeshie8 (ganeshie8):

write 10 as 1+9 and expand using binomial theorem

OpenStudy (anonymous):

oh, ok... (9 + 1)^999 - 1 = 9^999 + 9^998 +9^997 + ... + 9 then ?

ganeshie8 (ganeshie8):

\(9999\ldots (999 \text{times}) \\~\\ = \color{red}{10^{999}}-1\\~\\ =\color{red}{(1+9)^{999}}-1\\~\\ =\color{red}{1 + \dbinom{999}{1}*9+ \dbinom{999}{2}*9^2+ \cdots +\dbinom{999}{999}*9^{999}}-1 \)

ganeshie8 (ganeshie8):

the answer is \(2\) il let you figure out why :)

OpenStudy (anonymous):

but is this same like yours ? = 9^999 + 9^998 +9^997 + ... + 9

ganeshie8 (ganeshie8):

Looks you forgot the binomial coefficients

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

i know the formula of binomial newtown formula : (a + b)^n = a^n + a^(n-1)b + a^(n-2)b^2 + ... + b^n i supposed a = 9 and b = 1, that's why i get like above

ganeshie8 (ganeshie8):

your formula is wrong, check again

ganeshie8 (ganeshie8):

Here is the correct formula 

OpenStudy (anonymous):

my net conection is lagging, cant read the last your coment. but yeah, i forgot the coefficient there. thanks

OpenStudy (anonymous):

ok, got it. all terms is divisible by 9^2. :) Thank you very much @ganeshie8

ganeshie8 (ganeshie8):

right, but that doesn't prove "2" is the largest exponent

OpenStudy (anonymous):

what, .... you mean the answer is not n = 2 ???

ganeshie8 (ganeshie8):

answer is 2 but ur reason is wrong

OpenStudy (anonymous):

oh.... okay i want to show you how the to get that : (1 + 9)^999 - 1 = (999, 0) * 9^0 + (999 , 1) * 9 + (999 , 2) * 9^2 + .... + (999 , 998) * 9^998 + (999, 999) * 9^999 - 1 cancels 1's, get = (999 , 1) * 9 + (999 , 2) * 9^2 + .... + (999 , 998) * 9^998 + (999, 999) * 9^999 = (999 * 9) + (999 * 998) * 9^2 + .... + (999 * 9^998) + (9^999) faktor out all terms by take the GCD is 9^2 so, = 9^2 (111 + 999 * 998 + .... + 999 * 9^996 + 9^997) obvious 9^2 (111 + 999 * 998 + .... + 999 * 9^996 + 9^997) is divisible by 9^2 am i correct @ganeshie8 ?